A ball is dropped from a window and hits the ground with a speed of 25 ft/s. How high above the ground is the window?


1
Expert's answer
2021-10-24T18:18:06-0400

Given:

"v_0=25\:\rm ft/s"

"g=32\:\rm ft/s^2"


The maximum height

"h_{\max}=\frac{v_0^2}{2g}=\frac{25^2}{2*32}=9.8\:\rm ft"